LINESHAPE STUDIES OF ROTATIONAL TRANSITIONS OF OZONE AND NITRIC ACID

\underline{GEORGES WLODARCZAK}, BILAL BAKRI, JEAN-MARCEL COLMONT, JEAN DEMAISON, FRANCOIS ROHART , Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523, Université de Lille 1, F-59655 Villeneuve d’Ascq, France ; GABRIELE CAZZOLI, LUCA DORE, CRISTINA PUZZARINI , Laboratory of Millimetre-Wave Spectroscopy of Bologna, Dipartimento di Chimica ”G. Ciamician”, Università di Bologna, Via Selmi 2, I-40126 Bologna, Italy.

In the framework of a contract provided by the European Space Agency, linewidth measurements have been performed on selected rotational lines of ozone and nitric acid, mostly located between 300 and 350 GHz, in the B and C bands of the MASTER platform. Broadening parameters have been determined for O_{3} in collision with N_{2} and O_{2} at five different temperatures in the 190-300K range. For nitric acid, only air broadening coefficients were measured at 296K. The experiments have been performed at LMSB and PhLAM by using both different experimental set-ups and different analysis procedures. At PhLAM, the true lineshapes were observed with a video-type spectrometer and compared to Voigt and Galatry theoretical models. At LMSB, a Galatry model has been used to analyse the second derivative of the natural line profiles recorded with a frequency modulation spectrometer. Several lines of ozone and nitric acid were measured in both laboratories. The results of the intercomparison will be discussed.